Arrange in order of increasing energy the terms that may arise from the following configurations: 2p13p1, 3p13d1.
Answer #1

3p13d1>3p12p1

Quantum number rules

3p13d1 =3+3=6

3p12p1 =2+3=5

So, energy high is terms 3p13d1